I am to write an equation of the line containing the given point and parallel to the given line. What is the quation of the line with this given information
(5,7), x+2y=7

Found 2 solutions by lwsshak3, solver91311:
Answer by lwsshak3(11628)

I am to write an equation of the line containing the given point and parallel to the given line. What is the equation of the line with this given information
(5,7), x+2y=7
**
x+2y=7
change to standard form of a straight line: y=mx+b, m=slope, b=y-intercept
change given equation to standard form
2y=-x+7
y=-x/2+7/2
slope, m, of this equation is -1/2
second equation parallel to given equation would have the same slope, m=-1/2, but its
y-intercept, b, might be different
..
second equation:
y=-x/2+b
solving for b using given point (5,7)
7=-5/2+b
b=7+5/2=14/2+5/2=19/2
ans:
second equation: y=-x/2+19/2

Answer by solver91311(24713)



If you know one point and the slope of a line you can write an equation to represent that line using the point-slope form:



where are the coordinates of the given point and is the calculated slope.

Since you want a line parallel to the given line, use the fact that parallel lines have identical slopes, to wit:



Since your given equation is in standard form, namely , you can calculate the slope of the given line by

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism